OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through the given points. (1,4) (0,1)

OpenStudy (anonymous):

You simply use point slope formula: (y-y1)=m(x-x1)

OpenStudy (anonymous):

k can u please show me how to do it

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{x2-x1}\]

OpenStudy (anonymous):

-3 -1

OpenStudy (anonymous):

plug in your points for the slope: (1-(4))/(0-1) (-3)/-1=3 Plug in for one of your points: y-1=3(x-0) =y=3x+1

OpenStudy (anonymous):

thanks